malykykyk:

Which of the following best completes the proof showing that ΔWXZ ~ ΔXYZ? triangles WXZ and XYZ that share side XZ with right angle XZW, WZ equals 10, XZ equals 5, and YZ equals 2 and one half Since segment XZ is perpendicular to segment WY, angles WZX and XZY are both right angles and congruent. The proportion ________ shows the corresponding sides are proportional, so the triangles are similar by the SAS Similarity Postulate. 5 over 10 equals 5 over 2 and one half 5 over 10 equals 5 over 5 10 over 5 equals 5 over 2 and one half 10 over 5 equals 2 and one half over 5

malykykyk:

jhonyy9:

so i think from these choices just you need checked get the right ,true choice 5 over 10 equals 5 over 2 and one half --- this maybe true 5/10 = 5/2,5 so 1/2 = 2 5 over 10 equals 5 over 5 or this 5/10 = 5/5 so 1/2 = 1 10 over 5 equals 5 over 2 and one half 10 over 5 equals 2 and one half over 5 try the next two in this way and will get easy the correct answer @supie